# CDsampling: An R Package for Constrained D-Optimal Sampling in Paid Research Studies

### v3

@@ 4   Summary
> beta.noint = beta[2:(p+1)]> ## Estimating beta0 (intercept) compared to true> btemp0 = abs(beta.estimate[,1,]) #abs(beta0_est - beta0), beta0 = 0> mse.b0.mean=apply(btemp0,02,mean, na.rm = TRUE)> mse.b0.sd=apply(btemp0,2,sd) #sd of beta0 mse> ## Estimating beta1-beta4 compared to true> btemp=1/3*(beta.estimate[,2:(p+1),]-beta.noint)^2 #square of beta’s error> btemp=sqrt(apply(btemp,c(1,3),sum, na.rm = TRUE))> mse.b5all.mean=apply(btemp,2,mean, na.rm = TRUE) #apply mean over column> mse.b5all.sd=apply(btemp,2,sd, na.rm = TRUE)> ## separated mse sd for each beta parameter> btemp1 = sqrt((beta.estimate[,2,] - beta[2])^2)> mse.b1.mean=apply(btemp1,2,mean, na.rm = TRUE) #apply mean to columns, each method’s beta_1 mse> mse.b1.sd=apply(btemp1,2,sd, na.rm = TRUE) #sd of beta1 mse> btemp2 = sqrt((beta.estimate[,3,] - beta[3])^2)> mse.b2.mean=apply(btemp2,2,mean, na.rm = TRUE) #apply mean to columns, each method’s beta_1 mse> mse.b2.sd=apply(btemp2,2,sd, na.rm = TRUE) #sd of beta1 mse> btemp3 = sqrt((beta.estimate[,4,] - beta[4])^2)> mse.b3.mean=apply(btemp3,2,mean, na.rm = TRUE) #apply mean to columns, each method’s beta_1 mse> mse.b3.sd=apply(btemp3,2,sd, na.rm = TRUE) #sd of beta1 mse> output_simulation_rmse = as.data.frame(rbind(cbind(btemp0, category="beta0"),cbind(btemp, category="all_coef"), cbind(btemp1, category="beta1"), cbind(btemp2,category="beta21"), cbind(btemp3, category="beta22")))> library(data.table)> output_simulation_rmse.long=melt(setDT(output_simulation_rmse), id.vars=c("category"),variable.name=’Method’)> colnames(output_simulation_rmse.long)[3]=’RMSE’> colnames(output_simulation_rmse.long)[1]=’Coef’
@@
> library(ggplot2)> library(dplyr)> output_simulation_rmse.long %>ggplot(aes(x=Coef, y=as.numeric(RMSE), fill=factor(Method))) ++ geom_boxplot(alpha=0.7) + stat_summary(fun=mean, color = "black", position = position_dodge(0.75),+ geom = "point", shape = 18, size = 2, show.legend = FALSE)+ scale_fill_grey(start = 0, end = 1,+ name="Design Method") + xlab("Coefficient")+ylab("RMSE")+ theme_bw()++ facet_wrap(~Coef, scales=’free’)
S4   Template functions of constraint index set $I$

### v4

@@ 4  Summary
> beta.noint = beta[2:(p+1)]> ## Estimating beta0 (intercept) compared to true> btemp0 = abs(beta.estimate[,1,]) #abs(beta0_est - beta0), beta0 = 0> mse.b0.mean=apply(btemp0,02,mean, na.rm = TRUE)> mse.b0.sd=apply(btemp0,2,sd) #sd of beta0 mse> ## Estimating beta1-beta4 compared to true> btemp=1/3*(beta.estimate[,2:(p+1),]-beta.noint)ˆ2 #square of beta’s error> btemp=sqrt(apply(btemp,c(1,3),sum, na.rm = TRUE))> mse.b5all.mean=apply(btemp,2,mean, na.rm = TRUE) #apply mean over column> mse.b5all.sd=apply(btemp,2,sd, na.rm = TRUE)> ## separated mse sd for each beta parameter> btemp1 = sqrt((beta.estimate[,2,] - beta[2])ˆ2)> mse.b1.mean=apply(btemp1,2,mean, na.rm = TRUE)> mse.b1.sd=apply(btemp1,2,sd, na.rm = TRUE) #sd of mse> btemp2 = sqrt((beta.estimate[,3,] - beta[3])ˆ2)> mse.b2.mean=apply(btemp2,2,mean, na.rm = TRUE)> mse.b2.sd=apply(btemp2,2,sd, na.rm = TRUE) #sd of mse> btemp3 = sqrt((beta.estimate[,4,] - beta[4])ˆ2)> mse.b3.mean=apply(btemp3,2,mean, na.rm = TRUE)> mse.b3.sd=apply(btemp3,2,sd, na.rm = TRUE) #sd of mse> output_simulation_rmse = as.data.frame(rbind(cbind(btemp0, category="beta0"),cbind(btemp, category="all_coef"), cbind(btemp1, category="beta1"), cbind(btemp2,category="beta21"), cbind(btemp3, category="beta22")))> library(data.table)> output_simulation_rmse.long=melt(setDT(output_simulation_rmse), id.vars=c("category"),variable.name=’Method’)> colnames(output_simulation_rmse.long)[3]=’RMSE’> colnames(output_simulation_rmse.long)[1]=’Coef’
@@
> library(ggplot2)> library(dplyr)> output_simulation_rmse.long %>%+ ggplot(aes(x=Coef, y=as.numeric(RMSE), fill=factor(Method))) ++ geom_boxplot(alpha=0.7) ++ stat_summary(fun=mean, color = "black", position = position_dodge(0.75),+ geom = "point", shape = 18, size = 2, show.legend = FALSE)++ scale_fill_grey(start = 0, end = 1, name="Design Method") ++ xlab("Coefficient")+ylab("RMSE")+ theme_bw()++ facet_wrap(˜Coef, scales=’free’)
S4  Template functions of constraint index set II